OpenStudy (anonymous):

A note will pay $20,000 at maturity 10 years from now. How much should you be willing to pay for the note now if money is worth 5.2% compounded continuously?

OpenStudy (anonymous):

P(10) = X*e^(rt) 20,000 = X * e^(5.2%*10) X= $ 11,890